The probability that equipment used in a hospital lab will need recalibration is given in the following table. A service firm is willing to provide maintenance and provide any necessary calibrations for a fee of $625 per month. Recalibration costs $550 per time. Which approach would be most cost-effective, recalibration as needed or the service contract? (Round the answer of "Expected cost of recalibration" to nearest whole number. Omit the "$" sign in your response.)

Number of recalibrations 0 1 2 3 4
Probability of occurrence .10 .20 .25 .20 .25
Expected cost of recalibration $ __________

Pick one of the three below:

  • Service contract

  • Recalibration

  • Either approach costs the same

_____
Answer #1

expected number of recalibration =xP(x)=0*0.1+1*0.2+2*0.25+3*0.2+4*0.25=2.3

therefore expected cost of recalibration =550*2.3 =1265

since cost of Recalibration is higher than Service contract

therefore Service contract wold be most cost effective,